OpenStudy (anonymous):

solve the equation: 2-x=64. that 2 has an negative exponential.

OpenStudy (anonymous):

x = -62...?

OpenStudy (jiteshmeghwal9):

x=2-64=?

OpenStudy (anonymous):

-x=64-2 -x=62. x=-62